OpenStudy (anonymous):

Find the measure of angle BAC if the measure of angle BAD=68, the measure of angle BAC=2x-3 and the measure of angle CAD=4x+5

OpenStudy (anonymous):

set them equal to each other and solve for x

OpenStudy (anonymous):

how though thats what am trying to get out

OpenStudy (anonymous):

What do you mean how

OpenStudy (anonymous):

like how to you finish the anwers

OpenStudy (anonymous):

what would u get

OpenStudy (anonymous):

2x-3 = 4x +5 2x = -8 x = -4 we shouldn't be getting a negative

OpenStudy (anonymous):

oh 2x - 3 + 4x + 5 =68 6x+2=68 x=11

OpenStudy (anonymous):

90 -68 should normally be how you do it, or 180 -68

OpenStudy (anonymous):